Circular No. 7158

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ca IN NGC 3120 S. Woodings, R. Martin, and A. Williams report their discovery of an

apparent supernova (R ∼ 17) on a CCD image taken with the Perth-Lowell 0.61-m reflector on Apr. 27.55 UT in the course of the Perth Astronomy Research Group’s automated supernova search at Perth Observatory. A further observation on Apr. 28.47 also showed the new star at the same position. The last previous image of the field was taken on Apr. 2, with no object visible to mag ∼ 19. A. Verveer and J. Biggs provide the following position for SN 1999ca from an unfiltered CCD image obtained on Apr. 30.514 with the 0.25-m Mike Candy Telescope at Perth Observatory: α = 10^{h}05^{m}22^{s}.92 ± 0^{s}.02, δ = −34^{o}12′41′′.3 ± 0′′.2 (equinox 2000.0), which is ∼ 35′′ north of the center of NGC 3120; the magnitude was then 17.0.

SUPERNOVA 1999by IN NGC 2841 C. Gerardy and R. Fesen, Dartmouth College, report that optical CCD

spectra (range 450–800 nm, resolution 0.5 nm) of SN 1999by (cf. IAUC 7156), obtained with the McGraw-Hill 1.3-m telescope at MDM Observatory on May 2.2 UT, show it to be a type-Ia supernova, ∼ 10 days before maximum. Absorption features are observed at 485, 500, 546, 575, and 612 nm, with a strongly blue continuum.

CCD magnitude by D. Hanžl, Brno, Czech Republic (cf. IAUC 7157): May 4.068 UT, V = 14.00 ± 0.02. Visual magnitude estimates: May 3.822, 13.8 (M. Lehký, Hradec Králové, Czech Republic); 3.849, 13.6 (K. Hornoch, Lelekovice, Czech Republic); 3.89, 14.3 (R. J. Bouma, Groningen, The Netherlands).

SUPERNOVAE 1999ac, 1999ad, 1999an, 1999be CCD magnitudes, unfiltered unless otherwise noted: SN 1999ac in

NGC 6063, Apr. 8.691 UT, 16.0 (S. Yoshida and K. Kadota, Ageo City, Japan); 13.722, 15.8 (Yoshida and Kadota); 14.678, 16.0 (Yoshida and Kadota); 19.092, V = 15.91 (D. Hanžl, Brno, Czech Republic; comparison star GSC 945.845, V = 10.90, B − V = +0.678); 22.972, V = 16.1 ± 0.1 (L. Kiss, G. Fűrész, and B. Csák, Piszkéstető, Konkoly Observatory, Hungary). SN 1999ad in NGC 6063, Mar. 21.13, 14.7 (S. Moretti, Forlí, Italy). SN 1999an in IC 755, Mar. 13.657, 15.0 (Yoshida and Kadota); 19.962, V = 15.2 ± 0.2 (Kiss and Csák). SN 1999be in CGCG 089-013, Mar. 31.534, 16.5: (Yoshida and Kadota); Apr. 13.675, 15.7 (Yoshida and Kadota); 16.662, 15.6 (Yoshida and Kadota).

1999 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green